slope=m= y2-y1=55.9-59.5
x2-x1=1984-1964
m=-3.6= -0.18
20
y-intercept =b=+61
Can someone explain the steps to find +61 .I course understand all that is needed for finding the line of best fit. I just don't know where b=+61 comes into play.

Answer by Alan3354(69443)


slope=m= y2-y1=55.9-59.5
x2-x1=1984-1964
m=-3.6= -0.18
20
y-intercept =b=+61
Can someone explain the steps to find +61 .I course understand all that is needed for finding the line of best fit. I just don't know where b=+61 comes into play.
--------------
No idea. You typed it.